OpenStudy (anonymous):

Find an explicit rule for the nth term of a geometric sequence where the second and fifth terms are 40 and 5000, respectively.

OpenStudy (cwrw238):

2nd term = a^r 5th term = ar^4 where a = first term and r = common ratio you can find r by dividing 5th term by the 2nd and equating it to 5000/40

OpenStudy (cwrw238):

then find a by dividing 40 by r nth term = ar^(n-1)

OpenStudy (cwrw238):

ar^4 / ar = r^3 = 5000/40 can u continue?

OpenStudy (anonymous):

im trying to.

OpenStudy (anonymous):

an = 8 • 5n - 2 an = 8 • 5n + 3 an = 8 • 5n - 1 an = 8 • 5n + 1

OpenStudy (anonymous):

my answer choices

OpenStudy (cwrw238):

can you work ou what r is?

OpenStudy (anonymous):

i have no clue what to do

OpenStudy (cwrw238):

r^3 = 5000/40 = 125 r is cube root of 125 - use ur calculator for this

OpenStudy (anonymous):

5 cubed is 125